OpenStudy (anonymous):

A line with slope m through (4, 7) has equation y=mx+ _____ What is b?

OpenStudy (mathstudent55):

\(y = mx + b\) Substitute x and y with the coordinates of the point you were given, and solve for b.

OpenStudy (snowfire):

"b" is where the line hits the y axis, which is when x=0

OpenStudy (anonymous):

but you have two unknowns, m and b

OpenStudy (snowfire):

That's what confused me too, so I was hoping someone else would handle that.

OpenStudy (mathstudent55):

The y-intercept will be in terms of m since m is a variable. \(y = mx + b\) \(7 = m \cdot 4 + b\) \(b + 4m = 7\) \(b = 7 - 4m\)